Data concerning a recent period’s activity in the Prep Department, the first processing department in a company that uses process costing, appear below: Materials Conversion Equivalent units in ending work in process inventory 2,170 870 Cost per equivalent unit $ 15.56 $ 5.63 A total of 22,000 units were completed and transferred to the next processing department during the period. Required:

1. Compute the cost of ending work in process inventory for materials, conversion, and in total.

2. Compute the cost of the units completed and transferred out for materials, conversion, and in total.
Answer #1
1) Cost of ending work in process
Materials 2,170 * 15.56 = 33765.2
CC 870 * 5.63 = 4898.1
Cost of ending work in process 38663 answer
2) cost of units completed
Materials 22,000 * 15.56 = 342320
CC 22000 * 5.63 = 123860
cost of units completed 466180 answer